OpenStudy (anonymous):

Please help for a medal? Be quick! Explain how the Quotient of Powers was used to simplify this expression. 5 to the fourth power, over 25 = 52

OpenStudy (anonymous):

\[\frac{ 5^{1/4} }{ 25 } = 52\] Like this?

OpenStudy (anonymous):

Close, \[\frac{ 5^{4} }{ 25 } = 5^{2}\]

OpenStudy (anonymous):

\[\frac{ 5^{4} }{ 5^{2} } = 25\] Because you have the same base 5 on numerator and denominator, you can subtract their power. \[5^{4-2} = 25\] \[5^{2}= 25\] \[25 = 25\]

OpenStudy (anonymous):

Here are the options: By simplifying 25 to 52 to make both powers base five, and subtracting the exponents By simplifying 25 to 52 to make both powers base five, and adding the exponents By finding the quotient of the bases to be, one fifth and cancelling common factors By finding the quotient of the bases to be, one fifthand simplifying the expression Would the answer be A?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcomed :)

OpenStudy (anonymous):

Since you helped before twice I will give you a fan and a medal this time

OpenStudy (anonymous):

fan and medal for you too :)